OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet?

OpenStudy (anonymous):

 Volume = x*(2x)*(x-3)

OpenStudy (anonymous):

so is it 4x-3?

OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is cubic feet. What polynomial represents the height of the box in feet?

OpenStudy (anonymous):

\[x*2x=2x^2 \] \[2x^2(x-2)=2x^3-4x^2\]

OpenStudy (anonymous):

so is it 4x-3?

OpenStudy (anonymous):

Sorry , it should have been (x-3). Correct me.

OpenStudy (anonymous):

Im not getting it